OpenStudy (anonymous):

The 4th grade class is taking a field trip with 38 people. They use a total of 8 vehicles. Each car holds 4 passengers and each minivan holds 6 passengers. How many of each type of vehicle did they have?

OpenStudy (anonymous):

3 minivans, 5 cars 3 * 6 = 18 5 * 4 = 20 18+20=38

OpenStudy (anonymous):

Thank you!